# Reinforcement Twinning: from digital twins to model-based reinforcement learning

### 31 Jan 2024 v2

@@ 50       Policy switching and weight cloning based on the decision tree (Figure 3)
The idle policy continues its training through different mechanisms. No special considerations are needed if the idle policy is the model-based policy: this continues to be trained on the digital twin regardless of whether it is active on the real system or not. On the other hand, if the idle policy is model-free, the learning occurs indirectly through two mechanisms that establish a “demonstration” from the model-based to the model-free: (1) the storage of transitions in the memory buffer $\mathcal{D}_{\mathrm{L}}$ and (2) the continuous training of the value function Q. The off-policy nature of the reinforcement learning algorithm here implemented allows the agent to learn from policies that are exploratory or simply not ”optimal”, according to the current Q function. In this case, when the model-free policy is idle, and the Q function continues to be updated by observing the model-based policy in action on the real system, a sort of “demonstration” mechanism is established. The fact that the model-free could learn by the demonstration and eventually overtake the lead as a live policy is arguably the most significant contribution of this work.
@@ 5.1 Wind Turbine Control
Particularly critical is the transition between regions 2 and 3 (region 2.5 in Fig.4,) where gusts or lulls could result in shifting of operating conditions, leading to conflicting requirements between torque and pitch controllers. This region is typically handled via set point smoothing, mitigating transients when switching from one control logic to another (Bianchi et al., 2007; Abbas et al., 2020).
@@
$\hat{y}(t_{k})=(1-\alpha)y(t_{k})+\alpha\cdot y(t_{k-1})\,,$ (17)
with $\alpha$ being the low-pass filter coefficient computed as $\alpha=\exp(-2\pi f_{c}\Delta t)$, with $f_{c}$ the minimum cut-off frequency, $y$ the generic raw signal and $\hat{y}$ its low-pass filtered version.
We employ this filter for the generator rotational speed, $\omega_{g}$ and the measured wind speed, $u_{\infty}$. From the latter, it is possible to define the target rotational speed $\tilde{\omega}_{g}$ and extracted power $\tilde{P}$ from the reference curves in Figure 4. In Region 2, these two variables are linked by the operating conditions, while in Region 3, these are defined by mechanical constraints. In below-rated conditions, the power is theoretically maximized if the rotor operates at its optimal tip-speed-ratio $\lambda_{r,*}=\omega_{r}R/u_{\infty}$. Thus, one can define the reference generator speed relying on the measured wind speed as
@@
while the for model-free counterpart, we define the reward as
@@
where $\alpha_{1}$ and $\alpha_{2}$ ensure that the two components of the cost function have comparable orders of magnitude. The choice of including both the velocity and the power extraction is justified by the fact that these parameters are independent, in the sense that one could extract the same amount of power with different generator velocities by acting on the torque, possibly violating mechanical constraints.
@@
The wind energy community has put significant effort into the development of digital twins of onshore and offshore wind turbines, with applications ranging from load estimations (Pimenta et al., 2020; Branlard et al., 2023) to real-time condition monitoring (Olatunji et al., 2021; Fahim et al., 2022). In this work, we solely focus on the rotor’s aerodynamic performance and the twinning process seeks to identify a closure law for the power coefficient from data, starting from (25) and (26). In fact, the BEMT, commonly used to obtain the $C_{p}$ curve, assumes that the flow is steady, and this is questionable in the presence of large fluctuations in the wind speed (Leishman, 2002). Moreover, the coefficient can change over time as a result of turbine wear (e.g. due to blade erosion) or in the presence of extreme conditions (e.g. ice, dirt or bug buildup) Staffell and Green (2014). This is why any strategy based on a pre-computed $C_{p}$ curve is likely sub-optimal (Johnson, 2004, 2006) and the idea of continuous updates of this closure law from operational data has merit. Although this exercise is purely illustrative, the approach can be readily extended to more complex and multi-disciplinary model formulations.
@@
with the parameters $\bm{w}_{p}=[c_{1},c_{2},\dots c_{12}]$ to be identified. This model is known to be valid up to a maximum pitch excursion of $\beta=30$ deg. Therefore, we limit the maximum allowable pitch command to this value in our experiments.
@@ 5.2 Trajectory control of flapping wings micro air vehicles
Reproducing these performances requires advanced knowledge of the flapping wing’s aerodynamics in the low Reynolds number regime, as well as the fluid-structure interaction resulting from the wing’s flexibility. These remain significant challenges for advanced numerical simulations (Fei et al., 2019; Xue et al., 2023). The proposed test case shows how simplified models of the wing’s aerodynamics could be derived in real-time while an FWAVs seeks to achieve its control target. The considered problem is illustrated in Figure 5(a). A drone initially at $\bm{x}_{0}=[x_{0},z_{0}]$ m and $\bm{\dot{x}}_{0}=0$ m/s is requested to move to $\bm{\tilde{x}_{f}}=[\tilde{x}_{f},\tilde{z}_{f}]$ m as fast as possible, and then hover in that position, i.e. $\dot{\bm{\tilde{x}}}_{f}=[0,0]$, by continuously adapting its flapping wing motion. We challenge the controller by introducing gusty wind disturbances $\bm{u}_{\infty}$,
@@
$\bm{h}(\bm{x}_{k})=\frac{\bm{x}^{2}_{f}}{2}\mbox{exp}\biggl{[}-\frac{1}{2}\biggl{(}\frac{\bm{x}_{k}-\bm{x}_{f}}{\sigma}\biggr{)}^{2}\biggr{]}$ (32)
@@
We consider a drone with mass $m_{b}=3$ g and a constant flapping frequency of $f=20$ Hz. Given the initial position $\bm{x}_{o}=[0,0]$ m, we set the target position at $\bm{x}_{f}=[5,5]$ m. We assume that the manoeuvre should take less than $n_{t}=100$ flapping cycles, hence we consider an observation time of $T_{o}=5$ s. This defines the observation time for the twinning. We fix the pitching amplitude to $A_{\alpha}=45$ deg and let the flapping amplitude vary in the range $A_{\phi}\in[50;88]$ (deg). These values are comparable to what is observed in hummingbird’s flight, although the harmonic parameterization in (29) oversimplifies their actual wing dynamics (Kruyt et al., 2014). The stroke plane angle is bounded in the range $\beta_{w}\in[-30,30]$ (deg).
@@
The approach is based on Blade Element Momentum Theory (BEMT), which partitions the wing into infinitesimal elements (see Figure 6(a)). Each element produces an infinitesimal aerodynamic force that must be integrated along the span to get the total lift and drag. In the case of the smooth flapping motion (29), the forces due to the flapping motion dominate all the other physical phenomena (Lee et al. (2016)) so that the lift $L_{w}$ and drag $D_{w}$ expressions are
@@
The force coefficients $C_{L,w}$ and $C_{D,w}$ in equations 37 and 38 are modeled similarly to Lee et al. (2016) and Sane and Dickinson (2001):
@@ 6 Results
This section presents the performances of the RT algorithm in controlling and assimilating a wind turbine subject to time-varying wind speed (T1, 5.1), a flapping wing micro air vehicle (FWMAV) against head wind (T2, 5.2) and a cryogenic tank exposed to fluctuating thermal load (T3, 5.3). The hyper-parameters selected for each test case and the respective simulation configurations are listed in Table 1. We use five different random seeds for our simulations to extract performance statistics across different initial states of the optimizers. For both T1 and T2 we take the gradients ensemble over $N_{z}=$5 realizations, whereas in T3 we considered $N_{z}=$1, for the overall limited level of noise in the exogenous disturbance. In all studies, the first policy to be deployed onto the real system is the model-free one. Finally, for all test cases, we set $n_{S}=$1, as detailed in Section 4.3.
@@
T3 (Sec. 5.3) 4.32 $\cdot 10^{5}$ 30 20 1 1000 1000 30 5 4 20 1 0.5
6.1 Learning Performances
The learning performances for all the three cases are reported in Figure 10 and their numerical values tabulated in Table 2. The Figure shows the mean control or assimilation performances (first or second column, respectively) as a dashed line, and their relative standard deviation as a shaded area. The joined model-based and model-free control paradigm learns how to control effectively the system in a limited number of iterations for all test cases, showing good sample complexity. Indeed, the average control performance ($\mathcal{L}_{a}$) is in the order of $\mathcal{O}(10^{-1})$ with respect to the initial condition, performing thus up to ten times better the initial policy. The high standard deviation at the beginning of the training is associated with the initial exploration phase of the model-free side, and it consistently decreases with the iterations, as it converges to a specific region of the search space, as shown in Figures 9(c) and 9(d) whereas in Figure 9(a) it is roughly constant. Analysing the adjoint-based assimilation performances, Figures 9(b), 9(d) and 9(f) show that the methodology is able to tailor the digital twin on the current system specifics in five to ten iterations. At the end of the training process, it obtains an $\mathcal{L}_{p}$ of $\mathcal{O}(10^{-1})$, $\mathcal{O}(10^{-2})$ and $\mathcal{O}(10^{-4})$, normalized with the performances of the initial guesses of $\bm{w}_{p}$, for T1, T2 and T3, respectively. This, in turn, reflects on the model-based policy optimization (Step 5). In fact, as the digital twin predicts with greater precision the dynamics of the system for different actuations also the quality of the policy optimized “offline” increases. This also affects the policy switch (Step 6), as it becomes more accurate in choosing the best policy to be deployed on the real system. The effect of Step 6 is illustrated in Fig. 11.
@@ 6.1 Learning Performances
All experiments begin with the model-free policy driving the interactions with the real environment, i.e. as “live policy”, while the model-based is initially set as “idle”. Then, if the model-based performs better for $P_{\pi}$ iterations, it takes over as “live” policy following the decision tree of Step 6, illustrated in Figure 3. This is what happens at $n_{e}=7$ in T1, and $n_{e}=4$ for both T2 and T3. When this happens, the two paradigms, model-based and model-free, start to interact, as shown in Figure 11 especially appreciable in Figures 10(a) and 10(c). We observed that as soon as the model-based takes the lead of the interactions with the real system the model-free control performance registers a steeper improvement. Arguably, this has two causes: (1) the model-based acts as an “expert” for the model-free, providing high quality samples recorded in $\mathcal{D}_{\mathcal{L}}$ from which the model-free loop can learn; and (2) the critic gets better at predicting the value associated to the recorded transitions. Interestingly, in both T1 and T3 (Figure 10(a), 10(c)), the model-free learns a better policy imitating the model-based counterpart, effectively “surpassing the master” and being selected to be the “live” policy again. Comparing the performances of the model-based and model-free loops, In T1 the model based performs slightly better, in T2 the model-free performs roughly 50% better than the model-based and in T3 they converge to the same policy, thus we could not conclude about the general performance of the model-based and model-free policy optimization loops.
@@
6.2 Wind turbine Control
The digital twin performances, parametrized by the best set of weights $\bm{w}_{p}^{*}$, are shown in Figure 12. Figure 11(a) shows the average prediction versus the average system evolution as a function of time, while Figre 11(b) in the middle plots one versus the other. We conclude that the digital twin successfully tracks the dynamics of the real system. However, it shows a a systematic under-estimation of roughly $\Delta\bm{s}\lessapprox 1$ (rad / s) or 9.55 (RPM), especially localized at low rotational speeds. Whether this bias is inherently linked with the chosen parametrization or an insufficient number of samples requires further investigation. It is worth of notice that, while in T2 and T3 the coefficients to be identified represent the “ground truth”, in this test case there is not such a baseline, as the initial coefficients $\bm{w}_{p}^{0}$, derived in the work of Saint-Drenan et al. (2020), describe a power curve in steady wind conditions. As a consequence, the RT algorithm optimizes them to adapt to the dynamic a time-varying wind field as the system progresses. Therefore, the average assimilator loss $\mathcal{L}_{p}\approx\mathcal{O}(10^{-1})$ highlights the capability of the algorithm to adapt online, possibly taking into consideration changes in the operative conditions of the system itself, i.e. blade erosion and relative decrease in aerodynamic performances.
@@ 6.2 Wind turbine Control
The control performances are summarized in Figure 13. The controller is able to keep the system close to the desired set-points, Figures 12(a) and 12(b). This is especially true for what concerns the power, always kept very close to $\tilde{P}$, even in noticeably varying wind speeds (Figure 12(b)). On the other hand, it consistently keeps the generator speed, $\omega_{g}$, below-rated conditions, of approximately $\Delta\bm{s}\approx 8$ rad/s. This lower-than-rated rotational speed reflects on the actuated torque, Figure 12(d), which is always in the middle between the allowed action space. In addition, the very fast actuation rate on the torque, $\dot{\tau}$ = 15 kN, allows the controller to rapidly adapt in case of abrupt change in the wind speed, while fine tuning the actuation with the pitch angle, $\beta$. The latter follows consistently the wind speed, which is desirable. At higher wind speeds than rated, i.e. $u_{\infty}>11.4$ m/s, it consistently plays on the blades’ angle to reduce the aerodynamic torque to keep the system close to desired conditions. We note that this lower-than-rated $\omega_{g}$ strategy be mitigated, if desired, tuning the normalization coefficients $\alpha_{1}$ and $\alpha_{2}$, possibly increasing the penalty commonly done in multi-objective optimization problems.
@@
This decrease in $C_{p}$ is better appreciated in Figure 14(b), which shows the power coefficient trend against wind speed. This shows a peak, at rated conditions $u_{\infty}\approx 11$ (m/s) with a $C_{p}=0.42$, steadily decreasing for higher winds, as expected in common wind turbine control strategies.
Figure 14: Reduction of the power coefficient $C_{p}$ against wind speeds.
The model-based and model-free policies are compared in Figure 15. The two optimization loops converge to essentially the same parametrizations, leading to the same policy functions, as also supported by the very similar mean cost functional $\mathcal{L}_{a}$ listed in Table 2. The only (minor) difference appearing to be the threshold in the actuation of the torque. In both cases, when the error goes on the rotational speed goes negative, i.e. $e_{i}=\omega_{g}-\tilde{\omega}_{g}<0$, the controllers reduce the pitch angle (Figures 14(a),14(c)) to increase the aerodynamic efficiency and, consequently, the harvested aerodynamic torque, with the goal of increase the rotational speed to get closer to the set point. The opposite is true when the generator is spinning too fast: the controller increases the angle of attack of the blades to slow the generator down. The integral of the error over the past revolution, $\int e$, essentially decides on the slopes of the control surface. Interestingly, looking at the torque actuation shown in Figures 14(b), 14(d), it seems that this control actuation is almost insensitive to this control input, relying on the scalar error solely. This control surface is directly connected with the objective of keeping the power around the rated value of $P=5MW$. In fact, as the error becomes negative, the controller increases the torque, to keep the power at the end of the generator constant ($P_{g}=\omega_{g}\cdot\tau_{g}$). On the other hand, when the error increases, it reduces the resistive torque, relying on the pitch actuator to reduce the speed while keeping the power constant.
@@
Lastly, we conclude this analysis investigating the robustness of the found models, shown in Figure 15(b). Specifically, we run one-hundred simulations, on randomized initial conditions and sampled disturbances, possibly unseen during the training phase, using the found policies parametrizations $\bm{w}_{a,mb}^{*}$, $\bm{w}_{a,mf}^{*}$ and assimilation parameters $\bm{w}_{p}^{*}$. The control performances, Figure 15(a) show a multi-modal distribution. The stronger peak is located in a good performance region, close to $L_{a}\lessapprox\mathcal{O}(10^{2})$. This supports the good performances shown above. However, it appears that the model is not able to extrapolate the control strategies in unseen conditions outside of a given “confidence” region. This may be due to several factors. A first hypothesis is that it could be beneficial to use more wind fields time series (and relative realizations) to get a broader picture of the possible operative conditions, possibly increasing their diversity, or considering an higher number of realizations $N_{z}$. Another way of possible improvement may lie in the function approximators used for the policy function, currently just a linear combination of the inputs, which could possibly increase the descriptive capabilities of the model. Nevertheless, we considering the control task successfully achieved for the nominal operative conditions and leave these investigations for future works more centred on wind turbines control. The assimilations performances, on the other hand, are reliably on-track with respect what experienced during the learning phase. The digital twin tracks the real system with very small deviation, $\mathcal{L}_{p}<\mathcal{O}(0.5\cdot 10^{5})$ for most situations. However, this model may also be improved to take into account both the minor bias presented in Figure 12 and to be more flexible in adapting to new conditions.
Figure 16: Robustness analysis of the control policies (a) and assimilated model (b)
Figure 16: Robustness analysis of the control policies (a) and assimilated model (b)
6.3 Trajectory control of flapping wings micro air vehicles
In this test case, the initial weights $\bm{w}_{p}^{0}=[a,b.c]$, pertaining the lift and drag closure (equation (41) and (42)), are drawn from three random distributions, and a stopping criterion was defined to interrupt the $n_{G}$ iterations if the cost falls below a given threshold ($10^{-2}$). Figure 16(a) shows the assimilation performance of the digital twin in predicting the real-system behavior. The two show a small deviation. This can be better appreciated in Figure 16(b), which shows that the states predicted by the digital twin (y-coordinates) corresponds with good agreement with the true states (x-coordinate) experienced by the real environment, thus lying on the diagonal of the plot.

### 25 Feb 2024 v3

@@ 50       Policy switching and weight cloning based on the decision tree (Figure 3)
The idle policy continues its training through different mechanisms. No special considerations are needed if the idle policy is the model-based policy: this continues to be trained on the digital twin regardless of whether it is active on the real system or not. On the other hand, if the idle policy is model-free, the learning occurs indirectly through two mechanisms that establish a “demonstration” from the model-based to the model-free: (1) the storage of “expert” transitions in the memory buffer $\mathcal{D}_{\mathrm{L}}$ and (2) the continuous training of the value function Q. The off-policy nature of the reinforcement learning algorithm here implemented allows the agent to learn from policies that are exploratory or simply suboptimal according to the current Q function. In this case, when the model-free policy is idle, and the Q function continues to be updated by observing the model-based policy in action on the real system, a sort of “demonstration” mechanism is established. The fact that the model-free could learn by the demonstration and eventually overtake the lead as a live policy is arguably the most significant contribution of this work.
@@ 5.1 Wind Turbine Control
Particularly critical is the transition between regions 2 and 3 (region 2.5 in Fig.4,) where gusts or lulls could result in conflicting requirements between torque and pitch controllers. This region is typically handled via set point smoothing, mitigating transients when switching from one control logic to another (Bianchi et al., 2007; Abbas et al., 2020).
@@
$\hat{y}(t_{k})=(1-\hat{\alpha})y(t_{k})+\hat{\alpha}\cdot y(t_{k-1})\,,$ (17)
with $\alpha$ being the low-pass filter coefficient computed as $\hat{\alpha}=\exp(-2\pi f_{c}\Delta t)$, with $f_{c}$ the minimum cut-off frequency, $y$ the generic raw signal and $\hat{y}$ its low-pass filtered version.
We employ this filter for the generator rotational speed, $\omega_{g}$ and the measured wind speed, $u_{\infty}$. From the latter, it is possible to define the target rotational speed $\tilde{\omega}_{g}$ and extracted power $\tilde{P}$ from the reference curves in Figure 4. In Region 2, these two variables are linked by the operating conditions, while in Region 3, these are defined by mechanical constraints. In below-rated conditions, the power is maximized if the rotor operates at its optimal tip-speed-ratio $\lambda_{r,*}=\omega_{r}R/u_{\infty}$. Thus, one can define the reference generator speed relying on the measured wind speed as
@@
while for the model-free counterpart, we define the reward as
@@
In both equations, $\alpha_{1}$ and $\alpha_{2}$ ensure that the two components of the cost function have comparable orders of magnitude. The choice of including both the velocity and the power extraction is justified by the fact that these parameters are independent, in the sense that one could extract the same amount of power with different generator velocities by acting on the torque, possibly violating mechanical constraints.
@@
The wind energy community has put significant effort into the development of digital twins of onshore and offshore wind turbines, with applications ranging from load estimations (Pimenta et al., 2020; Branlard et al., 2023) to real-time condition monitoring (Olatunji et al., 2021; Fahim et al., 2022). In this work, we solely focus on the rotor’s aerodynamic performance and the twinning process seeks to identify a closure law for the power coefficient from data, starting from (25) and (26). In fact, the BEMT, commonly used to obtain the $C_{p}$ curve, assumes that the flow is steady, and this is questionable in the presence of large fluctuations in the wind speed (Leishman, 2002). Moreover, the coefficient can change over time as a result of turbine wear (e.g. due to blade erosion) or in the presence of extreme conditions (e.g. ice, dirt or bug buildup) (Staffell and Green, 2014). This is why any strategy based on a pre-computed $C_{p}$ curve is likely sub-optimal (Johnson, 2004, 2006) and the idea of continuous updates of this closure law from operational data has merit. Although this exercise is purely illustrative, the approach can be readily extended to more complex and multi-disciplinary model formulations.
@@
with the parameters $\bm{w}_{p}=[c_{1},c_{2},\dots c_{12}]$ to be identified. This model is known to be valid up to a maximum pitch excursion of $\beta=30$ deg. Therefore, we limit the maximum allowable pitch command to this value in our numerical experiments.
@@ 5.2 Trajectory control of flapping wings micro air vehicles
Reproducing these performances requires advanced knowledge of the flapping wing’s aerodynamics in the low Reynolds number regime, as well as the fluid-structure interaction resulting from the wing’s flexibility. These remain significant challenges for advanced numerical simulations (Fei et al., 2019; Xue et al., 2023). The proposed test case shows how simplified models of the wing’s aerodynamics could be derived in real-time while an FWAVs seeks to achieve its control target. The considered problem is illustrated in Figure 5(a). A drone initially at $\bm{x}_{0}=[x_{0},z_{0}]$ m and $\bm{\dot{x}}_{0}=0$ m/s is requested to move to $\bm{\tilde{x}_{f}}=[\tilde{x}_{f},\tilde{z}_{f}]$ m as fast as possible, and then hover in that position (i.e. $\dot{\bm{\tilde{x}}}_{f}=[0,0]$) by continuously adapting its flapping wing motion. We challenge the controller by introducing gusty wind disturbances $\bm{u}_{\infty}$,
@@
$\bm{h}(\bm{x}_{k})=\frac{\bm{x}^{2}_{f}}{2}\mbox{exp}\biggl{[}-\frac{1}{2}\biggl{(}\frac{\bm{x}_{k}-\bm{x}_{f}}{\sigma}\biggr{)}^{2}\biggr{]}\,.$ (32)
@@
We consider a drone with mass $m_{b}=3$ g and a constant flapping frequency of $f=20$ Hz. Given the initial position $\bm{x}_{o}=[0,0]$ m, we set the target position at $\bm{x}_{f}=[5,5]$ m. We assume that the manoeuvre should take less than $n_{t}=100$ flapping cycles, hence we consider an observation time of $T_{o}=5$ s. This defines the observation time for the twinning. We fix the pitching amplitude to $A_{\alpha}=45$ deg and let the flapping amplitude vary in the range $A_{\phi}\in[50;88]$ deg. These values are comparable to what is observed in hummingbird’s flight, although the harmonic parameterization in (29) oversimplifies their actual wing dynamics (Kruyt et al., 2014). The stroke plane angle is bounded in the range $\beta_{w}\in[-30,30]$ deg.
@@
The approach is based on BEMT, which partitions the wing into infinitesimal elements (see Figure 6(a)). Each element produces an infinitesimal aerodynamic force that must be integrated along the span to get the total lift and drag. In the case of the smooth flapping motion (29), the forces due to the flapping motion dominate all the other physical phenomena (Lee et al. (2016)) so that the lift $L_{w}$ and drag $D_{w}$ expressions are
@@
The force coefficients $C_{L,w}$ and $C_{D,w}$ in equations (37) and (38) are modeled similarly to Lee et al. (2016) and Sane and Dickinson (2001):
@@ 6 Results
This section presents the performances of the RT algorithm in controlling and assimilating a wind turbine subject to time-varying wind speed (T1, Sect. 5.1), a flapping wing micro air vehicle (FWMAV) against head wind (T2, Sect. 5.2) and a cryogenic tank exposed to fluctuating thermal load (T3, Sect. 5.3). The hyper-parameters selected for each test case are listed in Table 1. We use five different random seeds for our simulations to extract performance statistics across different initial states of the optimizers. For both T1 and T2 we take the gradients ensemble over $N_{z}=$5 realizations, whereas in T3 we considered $N_{z}=$1, given the lower level of noise in the exogenous disturbance. Finally, for all test cases, we set $n_{S}=$1, as mentioned in Section 4.3.
Section 6.1 presents the overall learning performances for both the twinning and the controller for all test cases, while the following sections report on the digital twins predictions and the identified control laws for each test case.
@@
T3 (Sect. 5.3) 4.32 $\cdot 10^{5}$ 30 20 1 1000 1000 30 5 4 20 1 0.5
@@
6.1 Learning Performances
The learning performances for all the three cases are reported in Figure 10 and their numerical values tabulated in Table 2. The Figure shows the mean control or assimilation performances (first or second column, respectively) with a dashed line, and a confidence interval with a standard deviation above and below the mean performances. The proposed RT paradigm learns to control effectively the system in a limited number of iterations for all test cases, showing good sample complexity. Indeed, the optimized policy performs up to ten times better than the initial one, as shown by the average control performance ($\mathcal{L}_{a}$). The large standard deviation at the beginning of the training of T2 and T3 can be explained by the initial exploration phase of the model-free loop. However, the standard deviation gets roughly constant as the policy optimization converges to a specific region of the search space, as shown in Figures 9(c) and 9(d). Analysing the adjoint-based assimilation performances, Figures 9(b), 9(d) and 9(f) show that the methodology is able to tailor the digital twin on the current system specifics within five to ten iterations. At the end of the training process, the assimilation reduces the initial cost $\mathcal{L}^{(0)}_{p}$ from the initial guess $\bm{w}_{p}$ by a factor $10$, $100$ and $10000$ for T1, T2 and T3, respectively.
This, in turn, reflects on the model-based policy optimization (Step 5). In fact, as the digital twin predictions of the system response improve, the quality of the policy optimized “offline” increases. This also affects the policy switch (Step 6), since the performances on the twin become more representative of the real responses. The effect of Step 6 is illustrated in Fig. 11. All experiments begin with the model-free policy driving the interactions with the real environment, i.e. as “live policy”, while the model-based is initially set as “idle”. As shown in Figure 11, the switch between model-free and model-based policies occurs at $n_{e}=7$ in T1 and $n_{e}=4$ for both T2 and T3, as $C_{\pi}>P_{\pi}$ in the policy switch in Figure 3. After the switch, the model-free control performance registers a steeper improvement despite remaining idle. This is due to a combination of factors: (1) the model-based acts as an “expert” for the model-free, providing high quality samples recorded in $\mathcal{D}_{\mathcal{L}}$ from which the model-free loop can learn; and (2) the critic function approximation gets better at predicting the value associated to the recorded transitions. Interestingly, in both T1 and T3 (Figure 10(a), 10(c)), the model-free learns a better policy by imitating the model-based counterpart, eventually “surpassing the master” and earning again the status of a “live” policy again within a few episodes.
Comparing the performances of the model-based and model-free loops, the model based performs slightly better than the model based in T1, but worst in T2 and equally in T3, where they converge to the same policy. The different performances in the different cases highlights the benefits of having two approaches to tackle the same problem and the merit in their combination.
@@ 6.1 Learning Performances
6.2 Performances on Wind Turbine Control
The digital twin performances on T1 are shown in Figure 12. Figure 11(a) shows the average prediction versus the average system evolution as a function of time, while Figure 11(b) plots one versus the other. We conclude that the digital twin successfully tracks the dynamics of the real system albeit with a systematic under-estimation of roughly 1 (rad /s) or 9.55 (RPM), especially at low rotational speeds. Whether this bias is inherently linked with the chosen parametrization or an insufficient number of samples requires further investigation. Differently from T2 and T3, the coefficients to be identified are unknown a priori, as the initial coefficients $\bm{w}_{p}^{0}$, derived in the work of Saint-Drenan et al. (2020), describe a power curve in steady wind conditions. Therefore, the average assimilation loss $\mathcal{L}_{p}\approx\mathcal{O}(10^{-1})$ highlights the capability of the algorithm to adapt online, possibly taking into consideration changes in the operative conditions of the system itself.
@@ 6.2 Performances on Wind Turbine Control
The control performances are summarized in Figure 13. The controller is able to keep the system close to the desired set-points, as shown in Figures 12(a) and 12(b). This is especially true for the power tracking, always kept very close to $\tilde{P}$ despite the noticeably varying wind speeds (Figure 12(b)). On the other hand, the controller consistently keeps the generator speed ($\omega_{g}$) below-rated conditions by approximately 8 rad/s. This lower-than-rated rotational speed reflects on the actuated torque, in Figure 12(d), which is always in the middle between the allowed action space. In addition, the very fast actuation rate on the torque ($\dot{\tau}$ = 15 kN/s) allows to rapidly adapt in case of abrupt changes in the wind speed, while fine tuning the actuation with the pitch angle $\beta$. The latter follows consistently the wind speed, which is desirable. At higher wind speeds than rated ($u_{\infty}>11.4$ m/s) it consistently plays on the blades’ angle to reduce the aerodynamic torque to keep the system close to desired conditions. This aerodynamic de-rating of the rotor is visible in Figure 14(b) which shows a peak of $C_{p}$ at rated conditions ($u_{\infty}\approx 11$ (m/s)) with a $C_{p}=0.42$, steadily decreasing for higher winds.
We note that the interplay between the two set-points on power and generator speed could be the modified fine-tuning the normalization coefficients $\alpha_{1}$ and $\alpha_{2}$, as commonly done in multi-objective optimization problems.
@@
Figure 14: Power coefficient $C_{p}$ against operative wind speeds.
The model-based and model-free policies are compared in Figure 15. The two optimization loops converge to essentially the same parametrizations, leading to the same policy functions, as also supported by the very similar mean cost functional $\mathcal{L}_{a}$ listed in Table 2. The only (minor) difference appears in the threshold of torque actuation. In both cases, when the error on the rotational speed goes negative ($e_{i}=\omega_{g}-\tilde{\omega}_{g}<0$), the controller reduces the pitch angle (Figures 14(a),14(c)) to increase the aerodynamic efficiency and, consequently, the aerodynamic torque, with the goal of increasing the rotational speed to get closer to the set point. The opposite is true when the generator is rotating too fast: the controller increases the angle of attack of the blades to slow the generator down. The integral of the error over the past revolution, $\int e$, essentially decides on the slopes of the control surface. Interestingly, looking at the torque actuation shown in Figures 14(b) and 14(d), it seems that this control actuation is almost insensitive to this control input, relying on the scalar error solely. This control surface is directly connected with the objective of keeping the power around the rated value of $P=5$ MW. In fact, as the error becomes negative, the controller increases the torque to keep the power at the end of the generator constant ($P_{g}=\omega_{g}\cdot\tau_{g}$). On the other hand, when the error increases, it reduces the resistive torque, relying on the pitch actuator to reduce the speed while keeping the power constant.
@@
We conclude this analysis by investigating the robustness of the identified models and the control laws. Figure 15(b) shows the distribution of the average (a) control cost function for both model based and model free and (b) assimilation, obtained by running one-hundred simulations on randomized initial conditions and wind profiles (unseen during the training phase) and using the identified model parameters $\bm{w}_{p}^{*}$ and policy parameters $\bm{w}_{a,mb}^{*}$, $\bm{w}_{a,mf}^{*}$.
The control performances shows a multi-modal distribution in Figure 15(b), with the strongest peak in the region of better performances ( $\mathcal{L}_{a}<\mathcal{O}(10^{2})$). This shows that at least one local minima exists within the range of investigate conditions and the training can eventually fall into the sub-optimal one depending on the combination of initial condition and wind profiles. A possible solution to mitigate the sensitivity to the training condition could be to extend the training to a broader range of wind profiles and/or initial conditions.
The assimilations performances, on the other hand, are reliably on-track with respect to what experienced during the learning phase. The digital twin tracks the real system with very small deviation ($\mathcal{L}_{p}<\mathcal{O}(0.5\cdot 10^{5})$) for most experienced operative conditions.
Figure 16: Robustness analysis of the assimilated closure (a) and control policies (b)
Figure 16: Robustness analysis of the assimilated closure (a) and control policies (b)
6.3 Performances on a FWMA Trajectory Control
In this test case, the initial weights $\bm{w}_{p}^{0}=[a,b,c]$ ((41) and (42)), are drawn from three random distributions, and a stopping criterion is defined to interrupt the $n_{G}$ iterations if the cost falls below a given threshold ($10^{-2}$). Figure 16(a) shows the assimilation performance of the digital twin in predicting the real-system behavior. The two show a small deviation. This can be better appreciated in Figure 16(b), which compares the real and predicted states.